40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Oct 17 23:28:32 2011, rana, Update, Electronics, StochMon 
    Reply  Wed Oct 19 18:18:03 2011, Suresh, Update, Electronics, StochMon : Filters installed PA190146.JPGPA190143.JPG
Message ID: 5706     Entry time: Wed Oct 19 18:18:03 2011     In reply to: 5682
Author: Suresh 
Type: Update 
Category: Electronics 
Subject: StochMon : Filters installed 

Quote:

To get to the bottom of the RFAM mystery, we've got to resurrect the StochMon to trend the RFAM after the IMC.

We will put an 1811 on the MC_TRANS or IP_POS beam (the 1811 has an input noise of 2.5 pW/rHz).

Then the Stochmon has an input pre-amp, some crappy filters, and then Wenzel RMS->DC converters. We will replace the hand-made filters with the following ones from Mini-Circuits which happen to match our modulation frequencies perfectly:

11 MHz     SBP-10.7+

55 MHz     SBP-60+

29.5 MHz   SBP-30+

 

The Stochmon had a four-way splitter, four hand-made filters and four mini-circuits ZX47-60-S+ Power Detectors.

Using the filters from our stock I have replaced the hand-made filters with the ones mentioned in Rana's elog.   The power supply solders to the ZX47-60-S+ Power Detectors were weak and came off during reassembly. And some of the handmade short SMA cables broke off at the neck.  So I changed the power supply cables and replaced all the short SMA cables with elbows.  I also removed one of the Power Detectors since there were four in the box and we need only three now.

The power supply connector on the box is illegal.  The current lab standard for  {+15, 0 , -15}  uses that connector.  So we are going to change it as soon as possible.  We need to identify a good {0, 5} lab standard and stock them.

The following were removed from the box:

PA190146.JPG

 

The box now looks like this:

PA190143.JPG

 

Steps remaining in installation of Stochmon:

1) Install the Newfocus 1811 PD at the IPPOS by diverting some of the power in that path

2) Connect the outputs of the Stochmon to ADC inputs in 1X2 rack.

ELOG V3.1.3-